OpenStudy (anonymous):

if tan^2(x) = sin^2(x)/cos^2(x) which by the half angle formulas = 1-cos(2x)/1+cos(2x)...the 1/2's cancel... does that meant that tan^2(x)=-1 or 0 or neither?

OpenStudy (turingtest):

neither, how do you figure it should be -1 or zero?

OpenStudy (anonymous):

\[1-\cos(2x)\div1+\cos(2x)\] take a -1 out of the top and you end up with \[-(-1+\cos(2x)\div1+\cos(2x)\]= -1? obviously this would be an identity if it were correct so i guess I'm just asking how it is wrong?

OpenStudy (turingtest):

\[-\frac{-1+\cos(2x)}{1+\cos(2x)}\neq-1\]check the algebra...

OpenStudy (turingtest):

\[\frac{1-\cos(2x)}{1+\cos(2x)}=-\frac{\cos(2x)-1}{1+\cos(2x)}\]but nothing is going to cancel here, so we just go round and round...

OpenStudy (anonymous):

ah I see. I think my lack of parenthesis made me see something that was not truly there. thank you.

OpenStudy (turingtest):

welcome :D